Use a cofunction to write an expression equal to csc(θ-45°)
Answer by Alan3354(69443)

Use a cofunction to write an expression equal to csc(t-45°)
-------
csc(t-45°) = 1/sin(t-45)
= 1/(sin(t)cos(45) - sin(45)cos(t))
= 1/(sqrt(2)/2)*(sin(t) - cos(t))
= sqrt(2)/(sin(t) - cos(t))
---------
Is that sufficient?